OpenStudy (anonymous):

Let f(x)=2 sqrt(x) If g(x) is the graph of f(x) shifted down 5 units and right 6 units, write a formula for g(x) g(x) =

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

I'm pretty sure this linear algebra, not calc. In order to solve this, you must first understand the basics of algebraic functions, which I think you do. To make this more understandable, lets first solve this visually, then algebraically. Let's re-write this function as an equation. \[f(x)=2\sqrt{x}\] \[y=2\sqrt{x}\] Now, lets graph this equation. It might look something like this (file attached).

OpenStudy (anonymous):

It is pre calc, transformation of functions is what my book has it under

OpenStudy (anonymous):

This is a square root function

OpenStudy (anonymous):

I'm not too clear on shifting vertically and horizontally and how that is represented in g(x)

OpenStudy (anonymous):

Ohhh alright so it's most likely brushing up on more advanced algebra.

OpenStudy (anonymous):

You're familiar with rise-run and point slope-form, correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

When shifting equations, it's sometimes easier to imagine the x and y axis visually and adjust accordingly. If: \[f(x)=2\sqrt{x}\] \[g(x)=2\sqrt{x-6}-5\]

OpenStudy (anonymous):

OK, that is what I was thinking... but I put that answer in and got it wrong

OpenStudy (anonymous):

If graphed:

OpenStudy (anonymous):

Thanks. That was it. So I had it like this: